# Graph Agent Network: Empowering Nodes with Decentralized Communications Capabilities for Adversarial Resilience

### v1

@@ III-E4 Decision Rule of $\mathcal{N}_{i}$
Then, according to the training samples $\mathcal{X}^{\mathcal{N}}=\mathcal{X}^{\mathcal{N}}_{1}\cup\mathcal{X}^{\mathcal{N}}_{0}$, the loss function of $\mathcal{D}_{i}$ is defined as the binary cross-entropy loss, i.e.,
$\small J_{\mathcal{N}}(\mathcal{X}^{\mathcal{N}})=-\frac{\sum_{x_{l}\in\mathcal{X}^{\mathcal{N}}_{1}}\log\left(\mathcal{N}_{i}(x_{l})\right)}{|\mathcal{X}^{\mathcal{N}}_{1}|}-\frac{\sum_{x_{l}\in\mathcal{X}^{\mathcal{N}}_{0}}\left(1-\log\left(\mathcal{N}_{i}(x_{l})\right)\right)}{|\mathcal{X}^{\mathcal{N}}_{0}|}.$ (11)
All loss functions are differentiable, enabling diverse functions to undergo self-training based on their associated losses. By employing SGD for parameter adjustment, these functions progressively converge the loss, thereby enhancing their inference capabilities.
III-F Action Update Rules
This section introduces action update rules, which play a crucial role in updating the actions of agents based on the actions (i.e., functions) received from their neighbors. By employing these rules, agents can dynamically adjust their actions to better adapt to the evolving environment and improve their collective performance.
In accordance with the previous discussion, we consider node $v_{i}$ as a case in point. From the perspective of an individual vertex $v_{i}$, it obtains the integrated feature of its neighbors’ actions in the current round by weighted fusion of the transmitted actions from its neighbors. Based on this synthesized feature, $v_{i}$ updates its own actions accordingly. The weighted fusion methods are elaborated in the following subsections.
III-F1 Action Update Rule of $\mathcal{M}_{i}$
The embedding function directly impacts the classification results which the adversaries target at, and any node in $\mathcal{V}$ could potentially be compromised (i.e., some neighbors of $v_{i}$ might be malicious). Therefore, $v_{i}$ cannot fully trust its neighbors’ embedding functions. Considering that the attention weights are individual functions, immune to the accumulation of malicious messages through iterative training, the weighted fusion of $\mathcal{M}_{i}^{\mathrm{rec}}$ using $\mathbf{w}_{i}$ can help filter out some harmful gradients, allowing for a benign update of $\mathcal{M}_{i}$. Specifically, the parameter of $\mathcal{M}_{i}$ is updated as
$\bm{\theta}_{\mathcal{M}_{i}}\leftarrow\bm{\theta}_{\mathcal{M}_{i}}+\eta_{\mathcal{M}}\frac{\omega_{i,j}\sum_{\mathcal{M}_{j}\in\mathcal{M}_{i}^{\mathrm{rec}}}\left(\bm{\theta}_{\mathcal{M}_{j}}-\bm{\theta}_{\mathcal{M}_{i}}\right)}{|\mathrm{N}_{i}|},$ (12)
@@ III-F1 Action Update Rule of $\mathcal{M}_{i}$
III-F2 Action Update Rule of $\mathcal{D}_{i}$
For the same feature, different degree inference functions may output different inference results. Therefore, $v_{i}$ aims to train a middleware function $\mathscr{D}_{i}$ such that the inference results for $\mathbf{z}_{k},\forall v_{k}\in\mathcal{V}$ derived from $\mathscr{D}_{i}$ are as similar as possible to those from recieved functions (i.e., the set $\mathcal{D}_{i}^{\mathrm{rec}}$). To address this, we randomly initialize $\mathscr{D}_{i}$ and design a loss function based on spatial random sampling to train $\mathscr{D}_{i}$, guiding the fusion of stored functions and the update of $\mathcal{D}_{i}$. Specifically, we perform the following steps: 1) Carry out small-scale random perturbations in the Euclidean space to sample features $\mathbf{z}_{s}$ similar to $\mathbf{z}_{i}$, which is calculated by $\mathbf{z}_{s}=\mathbf{z}_{i}+\xi\mathbf{x}$ where $\mathbf{x}$ is an i.i.d. 0-1 random vector that satisfies Gaussian distribution, $\xi$ is the sample range, thereby increasing the number of samples. 2) Utilize the features of nodes in $Z^{\mathrm{II}}_{i}$ as negative samples to enhance sample diversity from a limited perspective. 3) Employ $\mathscr{D}_{i}$ and $\mathcal{D}_{i}^{\mathrm{rec}}$ to infer these samples separately and use the mean squared error (MSE) to quantify the discrepancy between the inference results. The training strategy for $\mathscr{D}_{i}$ is illustrated in Fig. 4.
Figure 4: The training strategy for $\mathscr{D}_{i}$.
@@ III-F2 Action Update Rule of $\mathcal{D}_{i}$
$J_{\mathscr{D}}(v_{i})=\sum_{\mathcal{D}_{j}\in\mathcal{D}_{i}^{\mathrm{rec}}}\omega_{i,j}\Big{[}-\sum_{Q}{\mathbb{E}}_{\mathbf{z}_{s}\sim P\left(\mathbf{z}_{i}\right)}{\mathrm{MSE}\left(\mathscr{D}_{i}(\mathbf{z}_{s}),\mathcal{D}_{j}(\mathbf{z}_{s})\right)}\\ -\sum_{v_{k}\in Z^{\mathrm{II}}_{i}}{\mathrm{MSE}\left(\mathscr{D}_{i}(\mathbf{z}_{k}),\mathcal{D}_{j}(\mathbf{z}_{k})\right)}\Big{]},$ (13)
@@
where $P$ is a sampling distribution in the Euclidean space, $Q$ defines the number of sample, and $\mathrm{MSE}(\cdot)$ is the MSE function. After training that employing $J_{\mathscr{D}}(v_{i})$ as the loss, the parameter of $\mathcal{D}_{i}$ is updated as
$\bm{\theta}_{\mathcal{D}_{i}}\leftarrow\bm{\theta}_{\mathcal{D}_{i}}+\eta_{\mathcal{D}}\left(\bm{\theta}_{\mathcal{D}_{i}}-\bm{\theta}_{\mathscr{D}_{i}}\right),$ (14)
where $\bm{\theta}_{\mathscr{D}_{i}}$ is the parameter of the well-trained $\mathscr{D}_{i}$.
III-F3 Action Update Rule of $\mathcal{N}_{i}$
From Eqs. (9) and (10), it can be observed that under a limited view, the training set consistently encompasses $z_{i}$ when training $\mathcal{N}_{i}$ for any node $v_{i}$. This suggests that $\mathcal{N}_{i}$ can only determine whether a neighboring relationship exists between itself and a node with a given feature, rather than generalizing this inferential ability to arbitrary pairs of nodes. The action update mechanism presents an opportunity for such generalization. By receiving the neighboring functions $\mathcal{N}_{i}^{\mathrm{rec}}$ through the I/O interface, these functions enable the expansion of $\mathcal{N}_{i}$’s receptive field on a richer dataset, allowing $\mathcal{N}_{i}$ to learn the inferential capabilities of its neighbors.
Specifically, we perform the following steps: 1) Construct a training set using the same method as in Eqs. (9) and (10). 2) Both $\mathcal{N}_{i}$ and $\mathcal{N}_{i}^{\mathrm{rec}}$ are utilized to predict the output of the same training samples simultaneously. 3) For a specific neighbor $v_{k}$, the output of $\mathcal{N}_{k}$ serves as a metric to evaluate the similarity between the inferential ability of $\mathcal{N}_{i}$ and $\mathcal{N}_{k}$. If they are similar, it indicates that $\mathcal{N}_{i}$’s inferential ability on the sample aligns with $\mathcal{N}_{k}$, and substantial adjustments are not required. On the contrary, if they differ, a larger loss value is returned, compelling $\mathcal{N}_{i}$ to adjust its parameters to approach the corresponding $\mathcal{N}_{k}$. 4) Through iterative training, all received $\mathcal{N}_{i}^{\mathrm{rec}}$ are involved in evaluating $\mathcal{N}_{i}$, leading to the completion of function fusion.
@@ III-F3 Action Update Rule of $\mathcal{N}_{i}$
The evaluation result of $\mathcal{N}_{k}$ for $\mathcal{N}_{i}$ is quantified using the absolute values between the outputs. After all functions in $\mathcal{D}_{i}^{\mathrm{rec}}$ have provided their evaluations, the loss is calculated by weighted summation according to the aforementioned method, which in turn is used to compute the required gradient updates for a single round update of $\mathcal{N}_{i}$:
@@
$\small\bm{\theta}_{\mathcal{N}_{i}}\leftarrow\bm{\theta}_{\mathcal{N}_{i}}-\eta_{\mathcal{N}}\frac{\partial\sum_{x_{k}\in\mathcal{X}^{\mathcal{N}}_{0}\cup\mathcal{X}^{\mathcal{N}}_{1}}\sum_{\mathcal{N}_{j}\in\mathcal{N}_{i}^{\mathrm{rec}}}\omega_{i,j}\left|\mathcal{N}_{i}(x_{k})-\mathcal{N}_{j}(x_{k})\right|}{\mathrm{deg}_{i}|\mathcal{X}^{\mathcal{N}}_{0}\cup\mathcal{X}^{\mathcal{N}}_{1}|\partial\bm{\theta}_{\mathcal{N}_{i}}}.$ (15)
@@
III-G Filtering Adversarial Edges
@@
Specifically, the entire detection process is grounded in the functions trained from the agents’ first-person perspectives, with each agent acting both as a detector and a detectee. The detection procedure consists of four steps, illustrated by taking node $v_{i}$ as the detector and node $v_{x}$ as the detectee:
@@ IV Theoretical Analysis
In this section, we present the theoretical foundations for the design of inference function structures within agents and demonstrate their computational universality for the target tasks. Firstly, we establish the global equivalence between GAgN and GraphSAGE.
Theorem 1 (Equivalence).
@@
Denote $\mathbb{M}(\mathcal{G})=\{\mathbf{z}^{\mathbb{M}}_{i}=\mathcal{M}_{i}(\mathbf{z}_{i}):v_{i}\in\mathcal{V}\}$ as the global embedding result of all nodes, and $\texttt{SAGE}(\mathcal{G})=\{\mathbf{z}^{\texttt{S}}_{i},...\mathbf{z}^{\texttt{S}}_{|\mathcal{V}|}\}$ as the global classification result of GraphSAGE. For any SAGE there exists $\mathbb{M}$ that
$\mathbb{M}(\mathcal{G})=\texttt{SAGE}(\mathcal{G}),\forall\mathcal{G}\in G,$ (16)
where $G$ is the set of all attributed graphs.
Proof.

### v2

@@ 2.3 Decision Rules
Then, according to the training samples $\mathcal{X}^{\mathcal{N}}=\mathcal{X}^{\mathcal{N}}_{1}\cup\mathcal{X}^{\mathcal{N}}_{0}$, the loss function of $\mathcal{D}_{i}$ is defined as the binary cross-entropy loss (denoted as $\mathrm{CE}(\cdot)$), i.e., $J_{\mathcal{N}}(\mathcal{X}^{\mathcal{N}})=\mathrm{CE}(\mathcal{X}^{\mathcal{N}}_{1},\mathcal{X}^{\mathcal{N}}_{0})$
All loss functions are differentiable, allowing functions to self-train based on their losses. Using SGD for parameter adjustment, these functions reduce the loss and improve inference capabilities.
2.4 Action Update Rules
Action update rules are essential for updating agents’ actions (i.e., functions) based on functions received from their neighbors. Using these rules, agents can dynamically adjust their actions to adapt to the evolving environment and enhance collective performance.
Node $v_{i}$ acquires the integrated feature of its neighbors’ actions through weighted fusion of the transmitted actions in the current round. Based on these synthesized features, $v_{i}$ updates its actions accordingly. The weighted fusion methods are detailed in the following subsections.
Embedding Func.
These functions affects classification results, which adversaries may target, and any node in $\mathcal{V}$ could be compromised (some neighbors of $v_{i}$ might be malicious). Thus, $v_{i}$ cannot fully trust its neighbors’ $\mathcal{M}_{\cdot}$. Since attention weights are individual functions and resist malicious message accumulation, the weighted fusion of $\mathcal{M}_{i}^{\mathrm{rec}}$ using $\mathbf{w}_{i}$ can filter harmful gradients, enabling a safe update of $\mathcal{M}_{i}$. Specifically, the parameter of $\mathcal{M}_{i}$ is updated as
$\bm{\theta}_{\mathcal{M}_{i}}\leftarrow\bm{\theta}_{\mathcal{M}_{i}}+\eta_{\mathcal{M}}\omega_{i,j}\sum_{\mathcal{M}_{j}\in\mathcal{M}_{i}^{\mathrm{rec}}}\left(\bm{\theta}_{\mathcal{M}_{j}}-\bm{\theta}_{\mathcal{M}_{i}}\right)/|\mathrm{N}_{i}|,$ (7)
@@ 2.4 Action Update Rules
Degree Inference Func.
For identical features, different degree inference functions may yield varying results. To harmonize these, $v_{i}$ trains a middleware function $\mathscr{D}_{i}$ to ensure that inference results for $\mathbf{z}_{k},\forall v_{k}\in\mathcal{V}$ by $\mathscr{D}_{i}$ closely match those from the received set $\mathcal{D}_{i}^{\mathrm{rec}}$. This involves initializing $\mathscr{D}_{i}$ randomly and designing a loss function based on spatial random sampling for training, aiming to integrate stored functions and update $\mathcal{D}_{i}$. The training involves: 1) Implementing minor random perturbations in Euclidean space to sample features $\mathbf{z}_{s}$ akin to $\mathbf{z}_{i}$, calculated as $\mathbf{z}_{s}=\mathbf{z}_{i}+\xi\mathbf{x}$, where $\mathbf{x}$ is an i.i.d. 0-1 Gaussian vector and $\xi$ the sample range. 2) Using the features of nodes in $Z^{\mathrm{II}}_{i}$ as negative samples to broaden the sample diversity. 3) Applying $\mathscr{D}_{i}$ and $\mathcal{D}_{i}^{\mathrm{rec}}$ to separately infer these samples and utilizing the mean squared error (MSE) to evaluate the differences in inference outcomes. Training strategies for $\mathscr{D}_{i}$ are depicted in Figure 4.
@@
$J_{\mathscr{D}}(v_{i})=\sum_{\mathcal{D}_{j}\in\mathcal{D}_{i}^{\mathrm{rec}}}\omega_{i,j}\Big{[}-\sum_{Q}{\mathbb{E}}_{\mathbf{z}_{s}\sim P\left(\mathbf{z}_{i}\right)}{\mathrm{MSE}\left(\mathscr{D}_{i}(\mathbf{z}_{s}),\mathcal{D}_{j}(\mathbf{z}_{s})\right)}\\ -\sum\nolimits_{v_{k}\in Z^{\mathrm{II}}_{i}}{\mathrm{MSE}\left(\mathscr{D}_{i}(\mathbf{z}_{k}),\mathcal{D}_{j}(\mathbf{z}_{k})\right)}\Big{]},$ (8)
@@
where $P$ is a sampling distribution in the Euclidean space, $Q$ defines the number of sample. After training that employing $J_{\mathscr{D}}(v_{i})$ as the loss, the parameter of $\mathcal{D}_{i}$ is updated as $\bm{\theta}_{\mathcal{D}_{i}}\leftarrow\bm{\theta}_{\mathcal{D}_{i}}+\eta_{\mathcal{D}}\left(\bm{\theta}_{\mathcal{D}_{i}}-\bm{\theta}_{\mathscr{D}_{i}}\right).$
Figure 4: The training strategy for middleware function $\mathscr{D}_{i}$.
Neighboring Confidence Func.
From Eqs. (5) and (6), under a limited view, the training set always includes $z_{i}$ when training $\mathcal{N}_{i}$ for any node $v_{i}$. This means $\mathcal{N}_{i}$ can only detect neighboring relationships for itself and a node with a given feature. The action update mechanism allows for generalization. By receiving neighboring functions $\mathcal{N}_{i}^{\mathrm{rec}}$ via the I/O interface, $\mathcal{N}_{i}$ can expand its scope to a richer dataset, learning from neighbors’ inferential capabilities. 1) Construct a training set as per Eqs. (5) and (6). 2) Use both $\mathcal{N}_{i}$ and $\mathcal{N}_{i}^{\mathrm{rec}}$ to predict outputs of the same samples. 3) For a neighbor $v_{k}$, $\mathcal{N}_{k}$’s output measures the similarity in inferential ability with $\mathcal{N}_{i}$. Similarity implies alignment; otherwise, a higher loss prompts $\mathcal{N}_{i}$ to adjust towards $\mathcal{N}_{k}$. 4) Iterative training with all received $\mathcal{N}_{i}^{\mathrm{rec}}$ completes function fusion.
@@
The evaluation of $\mathcal{N}_{k}$ for $\mathcal{N}_{i}$ is quantified by the absolute differences between their outputs. After all functions in $\mathcal{D}_{i}^{\mathrm{rec}}$ provide evaluations, the loss is calculated by a weighted sum. This loss is then used to compute the gradient updates for a single update round of $\mathcal{N}_{i}$:
@@
$\bm{\theta}_{\mathcal{N}_{i}}\leftarrow\bm{\theta}_{\mathcal{N}_{i}}-\eta_{\mathcal{N}}\frac{\displaystyle\partial\left(\sum_{x_{k}\in\mathcal{X}^{\mathcal{N}}}\sum_{\mathcal{N}_{j}\in\mathcal{N}_{i}^{\mathrm{rec}}}\omega_{i,j}\left|\mathcal{N}_{i}(x_{k})-\mathcal{N}_{j}(x_{k})\right|\right)}{\partial\bm{\theta}_{\mathcal{N}_{i}}\mathrm{deg}_{i}|\mathcal{X}^{\mathcal{N}}_{0}\cup\mathcal{X}^{\mathcal{N}}_{1}|}.$ (9)
@@
2.5 Filtering Adversarial Edges
@@
Specifically, the entire detection process is grounded in the functions trained from the agents’ first-person perspectives, with each agent acting both as a detector and a detectee. The detection procedure is illustrated by taking node $v_{i}$ as the detector and node $v_{x}$ as the detectee:
  • 1.
@@ 2.6 Theoretical Analysis
This section provides the theoretical basis for designing inference function structures in agents, proving their computational universality for target tasks. We first establish the global equivalence between GAgN and GraphSAGE.
Theorem 1 (Equivalence):
@@
Denote $\mathbb{M}(\mathcal{G})=\{\mathbf{z}^{\mathbb{M}}_{i}=\mathcal{M}_{i}(\mathbf{z}_{i}):v_{i}\in\mathcal{V}\}$ as the embedding result of all nodes, $\texttt{SAGE}(\mathcal{G})=\{\mathbf{z}^{\texttt{S}}_{i},...\mathbf{z}^{\texttt{S}}_{|\mathcal{V}|}\}$ as the classification result of GraphSAGE, and $G$ as the set of all attributed graphs. For any SAGE there exists $\mathbb{M}$ that
$\mathbb{M}(\mathcal{G})=\texttt{SAGE}(\mathcal{G}),\forall\mathcal{G}\in G.$ (10)
Proof in Appendix A.1. We then propose the corollary to guide the design of the embedding function:
Corollary 1:
If model $\mathcal{M}_{i}$ is a single-layer neural network with trainable parameters as a $d_{z}\times d_{L}$ matrix, then $\mathbb{M}_{i}$ is universally capable of embedding any feature space onto the corresponding label space.
Proof in Appendix A.2. For the structure of the degree inference function, we arrive at the following conclusions:
Theorem 2:
Given an attribute graph $\forall\mathcal{G}\in G$, it is possible to map its feature matrix to any label matrix using only a single linear transformation (i.e., a trainable matrix) and a nonlinear activation.
Proof in Appendix A.3. For the neighboring confidence function’s structure, we present the following findings:
Theorem 3:
For any graph $\mathcal{G}$, there exists a fixed $2d_{z}$-dimensional weight vector $\mathbf{q}=[q_{1},\ldots,q_{2d_{z}}]$, such that for any two nodes $v_{i}$ and $v_{j}$ in $\mathcal{G}$ and their true neighboring relationship $\phi$ (where $\phi$ is a relative minimum or converges to 1), $\mathrm{CONCAT}_{r}(\mathbf{z}_{i},\mathbf{z}_{j})\mathbf{q}^{\top}=\phi$ is solvable.
Proof in Appendix A.4.
Table 1: Classification accuracy (%) under primary and secondary attackm with the highest scores in bold and the second highest underlined. $r_{p}$ is the perturbation rate. SVD and Pro is the abbreviation of GNN-SVD and Pro- GNN respectively.
$r_{p}$ RGCN SVD Pro Jaccard EGNN GAgN RGCN SVD Pro Jaccard EGNN GAgN
Cora 20 58.67 57.01 63.94 72.51 69.02 77.01 49.17±0.89 47.91±2.18 55.04±1.11 63.11±1.84 59.52±1.39 74.91±0.74
Citeseer 20 62.53 57.29 56.24 66.21 64.94 70.49 54.03±2.00 48.79±1.77 47.74±1.59 57.71±1.57 56.44±1.94 69.19±0.91
Polblogs 20 58.36 54.87 73.10 69.87 75.42 80.92 49.86±1.17 46.37±2.71 64.60±1.35 61.37±2.07 66.92±1.54 78.42±2.38
Pubmed 20 71.20 81.24 82.82 76.39 79.06 82.98 62.70±1.34 72.74±0.94 74.07±0.75 67.89±0.73 70.56±0.79 78.29±0.67
Figure 5: Attentions of random direction & inverse direction on the same edges.
Figure 7: Well-trained agents generated embedding.
3 Experiments
Datasets. Our approaches are evaluated on six real-world datasets widely used for studying graph adversarial attacks (Sun et al. 2020; Liu et al. 2022; Entezari et al. 2020; Li et al. 2022; Liu et al. 2024), including Cora, Citeseer, Polblogs, and Pubmed.
Baselines. The GAgN model protects non-defense GNNs from edge-perturbing attacks and outperforms other defensive approaches. Baseline models are evaluated as follows: Comparison defending models. GAgN is compared against: 1) RGCN, which uses Gaussian distributions for node representations to mitigate adversarial effects, 2) GNN-SVD, which applies truncated SVD for a low-rank adjacency matrix approximation, 3) Pro-GNN (Jin et al. 2020), which enhances GNN robustness through intrinsic node properties, 4) Jaccard (Wu et al. 2019), defending based on Jaccard similarity, 5) EGNN (Liu et al. 2021), filtering perturbations via $\ell_{1}$- and $\ell_{2}$-based smoothing. Attack methods. Experiments incorporate: 1) Metattack, a meta-learning based strategy, 2) G-EPA (Liu et al. 2022), a generalized edge-perturbing attack approach. The perturbation rate for attacks is set at 20%, a standard in the field unless otherwise specified.
Classification Accuracy
We evaluate the global classification accuracy of the proposed GAgN against primary and secondary attacks. Our experimental design uses GCN and its defenses as surrogates for Metattack and G-EPA, where Metattack represents the most potent primary attack and G-EPA an effective secondary strategy. Accuracy results are detailed in Table 1. As GAgN closes global input-output interfaces, it cannot directly serve as a surrogate in secondary attacks. To compensate, we transfer perturbed graphs from baseline models under secondary attacks to the GAgN test set. Five datasets are conducted, presenting average accuracies and variation ranges to illustrate GAgN’s resilience against secondary attacks. The diversity of the baselines likely covers potential vulnerabilities in defenses.
Our results indicate that GAgN achieves the best classification accuracy under both primary and secondary attacks. This is particularly noteworthy, as the lack of global input-output interfaces in GAgN makes it difficult for adversaries to exploit its vulnerabilities through global training, rendering it nearly impervious to secondary attacks.
Effectiveness of Aggregator
Symmetry of Attentions. In a well-trained GAgN, adjacent agents should learn similar attentions for their connecting edges (i.e., $\omega_{i,j}\approx\omega_{j,i}$) under the absence of collusion. To validate this, we first arbitrarily assign directions to the edges in the graph, defining the attention of any edge $e_{i,j}$ as $\omega_{i,j}$ from node $v_{i}$ to $v_{j}$. This results in a new graph with reversed attentions for all edges. We use a line chart to compare forward and reverse attentions for the same edges after learning in the GAgN model. For a certain edge’s smoothness, we sort the forward attention, obtain the sorted edge index, and use this index to find the corresponding reverse attention, examining variations in attentions. As shown in Figure 7, the variations between forward and reverse attentions are nearly identical, indicating that adjacent agents learn similar attentions. This confirms the effectiveness of the neighbor feature aggregator in the GAgN model. To demonstrate consistency among adjacent agents in learning attention for the same edge, we selected the node with the highest degree (168) from Cora. Experimental results, presented in Appendix B.1, show the learning curves of neighboring agents’ attentions on the same edge.
Distribution of Attentions. Here we investigate the attention distribution learned by agents for both normal edges and adversarial edges. A significant distinction between the two would indicate that agents, through communication, have become aware that lower weights should be assigned to adversarial edges, thereby autonomously filtering out information from illegitimate neighbors introduced by such edges. As the previous set of experiments have demonstrated the symmetry of attention, we present the attention of a randomly selected agent on one end of the edge. The experimental results, as depicted in Figure 7, reveal a notable difference in the kernel density estimation (KDE) of attentions between normal edges and adversarial edges (generated by Metattack). This outcome substantiates that agents, by training the aggregator, have acquired the capability to autonomously filter out adversarial edges.
Effectiveness of Embedding
After communication, the agent uses a well-trained embedding function to map its feature into the label space. We visualize the global embeddings to assess their effectiveness. If agents’ embeddings, trained only on their labels under limited knowledge, still show global clustering, it indicates the embedding function’s success. Figure 7 shows the t-SNE visualization of Cora and Citeseer embeddings (Van der Maaten and Hinton 2008), with points of different colors indicating distinct categories. The clear clustering of points validates the embedding function’s efficacy. Additionally, the visualization of the embedding process is presented in Appendix B.2, illustrating how GAgN progressively learns effective embeddings.
Table 2: Degree inference accuracy (%)
Cora Citeseer Polblogs Pubmed
L-N 90.35 89.21 92.41 81.08
S-N 95.31 94.19 96.32 83.45
D-N 92.64 91.08 92.57 83.06
Effectiveness of Degree Inference
Each agent possesses the ability to infer the degree of any node based on the degree inference function. To validate the effectiveness of this inference, we select 150 representative nodes as test sets after training the GAgN model on the corresponding clean graph, following these rules: 1) the top 50 nodes with the highest degree, denoted as L-N; 2) the top 50 nodes with the lowest degree, denoted as S-N; 3) the 50 nodes furthest away from the inference node in terms of graph distance, denoted as D-N. Notably, to avoid training set leakage into the test set, an agent’s 1-hop neighbors and select 2-hop neighbors are excluded from its test set. The experimental results, as shown in Table 3, demonstrate that even without data on the test nodes, agents can generalize their inference capabilities to the nodes of the graph under limited knowledge training, effectively inferring the degree of nodes.
Effectiveness of Neighboring Confidence
To validate the neighboring confidence function on graph $\mathcal{G}$, $|\mathcal{V}\times\mathcal{V}|$ matrix operations are needed. To reduce computational complexity and ensure fair evaluation, we construct four representative graphs and use their neighbor relationships as test samples: 1) The original graph $\mathcal{G}_{0}$, with all labels set to 1. 2) A graph $\mathcal{G}_{1}$ with randomly distributed edges, labeled based on whether edges are original. 3) A randomly rewired graph $\mathcal{G}_{2}$, with all labels set to 0. 4) A graph $\mathcal{G}_{3}$ with only adversarial edges, all labeled 0. We randomly select 50 nodes to compute the confidence for neighbor relationships, treating this as a binary classification problem. Classification accuracy measures the function’s effectiveness. Figure 8 shows that the neighboring confidence function performs well across all test sets, demonstrating its effectiveness.
Figure 8: Accuracy of neighboring Confidence function.
Figure 9: ROC curves with different attention threshold.
Ablation Experiments
We employed a multi-module filtering approach to reduce computational complexity when filtering adversarial edges. To investigate the individual effects of these modules when operating independently, it is crucial to conduct ablation experiments. Section 3 has explored the distribution patterns of attention scores for normal and adversarial edges. Edges with low attention scores are initially considered suspicious, which implies that the attention threshold may influence the detection rate and false alarm rate. To quantitatively examine the impact of threshold selection on these rates, we utilize receiver operating characteristic (ROC) curves. As depicted in Figure 9, the detection rate is relatively insensitive to the attention threshold. This allows for a high detection rate while maintaining low false alarm rates.
Table 3: Results of ablation experiments
Dataset Cora Citeseer
Metrics TP (TPR) FP (FPR) TP (TPR) FP (FPR)
Att. 928 (85.52%) 302 (5.56%) 782 (82.66%) 341 (7.20%)
Att.+$\mathcal{D}$+$\mathcal{N}$ 901 (83.04%) 47 (0.87%) 753 (79.60%) 55 (1.16%)
After identifying suspicious edges based on attention, the degree inference and neighboring confidence functions help filter out normal edges, reducing false positives. We examine the effectiveness of these functions in detecting adversarial edges while retaining normal ones. True positives (TP) and true positive rate (TPR) measure detection accuracy, while false positives (FP) and false positive rate (FPR) measure normal edge misclassification. We use Metattack for testing. Table 3 shows results: “Att.” refers to detection by attention alone, and “Att.+$\mathcal{D}$+$\mathcal{N}$” refers to combined detection. Detection using only attention has a higher rate but also more false positives. Adding degree inference and neighboring confidence reduces false positives significantly without majorly affecting detection rates.
4 Conclusions
Appendix A Proofs
A.1 Proof of Theorem 1